Find the work possess for a Helium gas at 200C.


Work = ____ kJ/kg


1
Expert's answer
2021-09-10T11:15:27-0400

Gives

"\\T=273+20=293K\\W=?"

Gas eqution

PV=nRT

"W=PV"

W=nRT

"W=\frac{m}{M}RT"


"W=\frac{1}{4}\times8.314\times(273+20)=609KJ/kg"